### v1

# MARS: Audio Generation via Multi-Channel Autoregression on Spectrograms

Abstract
Research on audio generation has progressively shifted from waveform-based approaches to spectrogram-based methods, which more naturally capture harmonic and temporal structures. At the same time, advances in image synthesis have shown that autoregression across scales, rather than tokens, improves coherence and detail. Building on these ideas, we introduce MARS (Multi-channel AutoRegression on Spectrograms), a framework that treats spectrograms as multi-channel images and employs channel multiplexing (CMX), a reshaping technique that lowers height and width without discarding information. A shared tokenizer provides consistent discrete representations across scales, enabling a transformer-based autoregressor to refine spectrograms from coarse to fine resolutions efficiently. Experiments on a large-scale dataset demonstrate that MARS performs comparably or better than state-of-the-art baselines across multiple evaluation metrics, establishing an efficient and scalable paradigm for high-fidelity audio generation.
1 Introduction
In recent years, audio generation has undergone significant advancements, primarily fueled by the rapid development of deep generative models. These methods can broadly be divided into two categories: those operating in the time domain and those working in the frequency domain.
Time-domain models attempt to synthesize audio waveforms directly, learning to capture both short- and long-term dependencies in the raw signal. A wide spectrum of architectures has been explored in this direction, including Variational Autoencoder (VAE)-based models [1], Generative Adversarial Network (GAN)-based approaches [2, 3], and denoising diffusion probabilistic models (DDPMs) [4]. While promising, such models often encounter difficulties in modeling the complex hierarchical structure of audio over longer time horizons, resulting in challenges related to scalability and perceptual fidelity, as underlined in [5].
An alternative line of research focuses on the frequency domain, where audio is represented as a spectrogram. In this case, generative models operate on two-dimensional time–frequency representations, which are later transformed back into waveforms via inverse short-time Fourier transform (ISTFT). Examples include GANSynth [6], which applies GANs to spectrogram generation, and EDMSound [7], which leverages diffusion models. This approach benefits from the fact that spectrograms explicitly capture harmonic structures and temporal envelopes, making it easier for models to learn recurring patterns. However, frequency-domain methods shift the challenge to the accurate reconstruction of fine-grained spectral details and phase information, both of which are critical for perceptually convincing audio.
In parallel, the domain of image generation has witnessed transformative progress. Traditional autoregressive models, which predict one token at a time in a sequential manner, have been surpassed by a new family of approaches known as Visual AutoRegressive (VAR) models [8]. VAR models redefine the autoregressive paradigm by predicting the next resolution scale rather than the next token. Starting from a coarse representation, the model progressively refines the output to higher resolutions, effectively capturing global structure at the lower scales and local details at finer scales. This next-scale prediction strategy significantly reduces sequence length, accelerates inference, and improves quality. Building upon this idea, subsequent refinements such as ImageFolder [9] have introduced stronger tokenizers and improved training objectives, further boosting performance in high-resolution image synthesis. By capturing both semantic consistency and fine-grained textures, VAR-based approaches have set a new state-of-the-art in image generation, with applications extending to domains where structural coherence is crucial.
Motivated by these developments, in this work we propose MARS (Multi-channel AutoRegression on Spectrograms), an autoregressive framework that adapts next-scale prediction to the audio domain. MARS treats spectrograms as multi-channel images and leverages a shared tokenizer to learn consistent discrete representations across scales. The autoregressive model then progressively predicts higher-resolution tokens conditioned on coarser ones, enabling hierarchical refinement of the generated spectrogram.
Fig. 1: Audio preprocessing pipeline for tokenizer input preparation and channel multiplexing (CMX) for reducing input resolution.
A central component of our approach is a novel preprocessing strategy called channel multiplexing (CMX), which reduces spatial resolution while redistributing information along the channel dimension. This technique allows us to preserve spectral fidelity while keeping computational cost manageable, making the method scalable to long and wide bandwidth audio recordings. By decoupling resolution from information density, CMX ensures that MARS can exploit the inductive biases of modern neural architectures optimized for multi-channel data.
To rigorously evaluate MARS, we rely on the NSynth dataset [10], a widely used benchmark for audio generation. Following the protocol of [11], we employ sample diversity metrics including NDB/$k$ and Inception-based scores for pitch and instrument classes, which assess whether the samples capture the distribution of musical attributes, and embedding-based similarity metrics such as kernel Inception distance (KID) and Fréchet audio distance (FAD), which compare reconstructed and reference audio distributions in learned feature spaces.
Our results demonstrate that MARS achieves competitive performance compared to leading models such as DDSP [12], DiffWave [4], and NSynth [10]. Beyond strong quantitative results, MARS also provides a new perspective on autoregressive modeling for audio, highlighting that thanks to its scale-wise refinement and channel multiplexing design, it achieves this level of quality while keeping computational costs contained, offering a favorable balance between performance and efficiency.
2 Proposed Method
Training AR models consists of two stages: first, learning a tokenizer that remains consistent across all resolutions; and second, training the AR model to predict higher-resolution tokens conditioned on the lower-resolution token map.
Fig. 2: Tokenizer architecture. The tokenizer is adapted from ImageFolder [9], improving upon the original VAR tokenizer [8]. The input spectrogram is partitioned into patches of size $L\times L$ and concatenated with $S$ learnable tokens before being processed by a transformer encoder $\mathcal{E}$, producing latent representations $z$. These are discretized by a vector quantizer to obtain $z^{\prime}$, which are then combined with another set of $L\times L$ learnable tokens and passed to a decoder $\mathcal{D}$ for reconstruction.
2.1 Audio Preprocessing
We first convert each waveform into a spectrogram using a STFT, retaining only the amplitude since the phase is later (after inference) reconstructed with the Griffin–Lim algorithm [13]. As an example, an 8s waveform at 16kHz with a 1024-point STFT and hop size of 256 yields a spectrogram of $512\times 512$, already matching the largest image size attempted with VAR [8]. Higher sampling rates further enlarge spectrograms, leading to excessive memory consumption. Moreover, processing such large inputs requires greater network depth to extend the receptive field, as convolutional layers scale only gradually; in fact, increasing input resolution from $256\times 256$ to $512\times 512$ in the original VAR model raised the parameter count from 310M to 2.3B.
To address this problem, which makes our experiments quickly unfeasible, we propose channel multiplexing (CMX), which reduces spatial dimensions by redistributing values across channels in a chessboard-like scheme (Figure 1). For spectrograms derived from an $N$-point STFT, the reshaping procedure jointly reorganizes the frequency and temporal dimensions, producing a tensor with the desired reduced spatial resolution and with a channel cardinality $C$ scaling with $N$, the signal length, and sampling rate. In this way, CMX accommodates inputs of varying length and sampling rate by adjusting $C$, while preserving the original data information and substantially reducing spatial overhead. As shown in Section 3.4, CMX maintains performance, supports larger batch sizes with cleaner gradients, and allows the use of architectures designed for smaller inputs—without the need to increase network depth, which is typically required to extend receptive fields—thereby avoiding significant parameter growth while enabling efficient processing of large-scale data. In addition, for our tokenizer presented in Figure 2, even without increasing network depth, enlarging the input dimension $M\times M$ leads to a parameter growth of $O(K^{2}+S)$ where $K=M/L$, $L$ is the patch size, and $S$ the number of learnable tokens, already scaling with the square of the input dimension.
More generally, CMX decouples information density from time–frequency resolution. Rather than discarding or compressing data, it reorganizes it into a format naturally suited to architectures optimized for multi-channel inputs (e.g., RGB images). This enables long recordings or high-fidelity signals to be processed with bounded memory cost, and the same principle can be extended beyond audio to images, video, medical scans, or structured time series. CMX thus provides a compact, lossless representation that scales efficiently, reducing memory and compute requirements while maintaining full data fidelity.
2.2 Tokenizer
The tokenizer architecture and training procedure are adapted from ImageFolder [9], extending the original VAR tokenizer [8]. The overall design is illustrated in Figure 2.
The tokenizer is trained using a composite objective that combines three terms:
$\mathcal{L}=\lambda_{recon}\mathcal{L}_{recon}+\lambda_{VQ}\mathcal{L}_{VQ}+\lambda_{ad}\mathcal{L}_{ad}.$ (1)
Here, $\mathcal{L}_{recon}$ is an $L_{2}$ reconstruction loss that ensures fidelity between the reconstructed and ground-truth spectrograms; $\mathcal{L}_{VQ}$ is the vector quantization loss, which aligns the encoder outputs with the nearest codebook entries; and $\mathcal{L}_{ad}$ is an adversarial loss applied through a PatchGAN discriminator [14], encouraging reconstructions to be indistinguishable from real spectrograms.
This design ensures that the tokenizer produces discrete representations that are both semantically meaningful and perceptually aligned, which is crucial for autoregressive refinement across scales.
Using a shared tokenizer across resolutions provides a crucial advantage, particularly for audio spectrograms. This is because sound often contains harmonics, which are higher-frequency components that are multiples of the fundamental frequency. In spectrograms, these harmonics reproduce the same temporal patterns as the fundamental frequency, but at finer frequency resolutions. A tokenizer that is consistent across scales can therefore encode these recurring structures in a unified way. As a result, this inductive bias allows the AR model to more effectively capture and reproduce these repeating structures, leading to more coherent and high-fidelity audio generation.
2.3 Autoregressive Model
Once the tokenizer is trained, the autoregressive model is trained following [8] to conduct the next-scale AR modeling for a faster inference speed.
The AR model is designed as a transformer-based architecture that progressively predicts tokens across different scales, effectively reducing the autoregressive sequence length and accelerating generation. During training, the model learns to condition predictions of fine-scale tokens on coarser ones, enabling a hierarchical refinement of the output. This multi-scale training strategy not only improves efficiency but also enhances consistency across scales, leading to higher-quality generated samples and reduced inference cost compared to standard autoregressive approaches.
Table 1: Reconstruction results according to the objective metrics reported in [11]. The best results are highlighted in bold, and the second best are underlined. The last (grayed) row reports metrics on generated audio (i.e., after the autoregressive module), and values are not directly comparable against those in the previous four rows.
Model NDB/$k$ ($\downarrow$) PKID ($\downarrow$) IKID ($\downarrow$) PIS ($\uparrow$) IIS ($\uparrow$) MSE ($\downarrow$) MAE ($\downarrow$) FAD ($\downarrow$)
Diffwave 0.74 0.0093 0.0021 2.3814 5.6477 0.0291 0.1369 7.9488
DDSP 0.20 0.0053 0.0020 3.3224 5.3371 0.0130 0.0666 1.1519
NSynth 0.74 0.0101 0.0024 2.3238 4.6364 0.0329 0.1224 4.0590
MARS (ours) 0.19 0.0035 0.0015 2.9602 5.2047 0.0143 0.0915 1.6429
[…]
3 Experimental Evaluation
3.1 Experimental Setup
We tested our autoregressive model using NSynth [10], which is a dataset of over 300,000 musical notes, each with a unique pitch, timbre, and envelope. The samples are mono signals of approximately 4s and with a sampling rate of 16 kHz. We converted them into spectrograms using a 1024-point STFT and rearranged them using CMX in tensors of dimension $256\times 256\times 2$, splitting frequencies among two channels. We trained our tokenizer for 400 epochs following the training settings proposed in [9] and our AR model for 350 epochs following [8]. We run our experiments on a single RTX 5000 Ada GPU with 32GB of RAM.
3.2 Evaluation Metrics
To assess our model, we follow the comprehensive evaluation protocol proposed by [11], which benchmarks state-of-the-art audio generation methods on the NSynth dataset. In line with this methodology, we evaluate our tokenizer for audio reconstruction using a diverse set of metrics that capture complementary aspects of audio. These metrics are grouped into three categories:
  • Reconstruction errors. We report MSE and MAE, computed on mel-spectrograms extracted from audio samples, as recommended in [11].
  • Sample diversity metrics. The Number of Statistically Different Bins (NDB/$k$) [15] measures sample diversity by clustering the training space and comparing the distribution of test samples to training data. Scores near 0 indicate good coverage of the training distribution, while values close to 1 reveal mode collapse. To further assess diversity and semantic consistency, we also adopt the Pitch Inception Score (PIS) and Instrument Inception Score (IIS) [16], which evaluate whether the reconstructed samples capture the true distribution of pitch classes and instrument classes in NSynth. Following [11], we also report these metrics, although they focus on “symbolic” properties of notes rather than audio quality.
  • Embedding-based metrics. Kernel Inception Distance (KID) [17] measures the similarity between the distributions of reference and reconstructed embeddings. Specifically, we compute PKID and IKID [16] using the same Inception networks employed for PIS and IIS, where the maximum mean discrepancy (MMD) is applied to compare the embedding distributions. We also adopt Fréchet Audio Distance (FAD) [18], a widely adopted metric for perceptual quality, which leverages VGGish embeddings of reference and reconstructed audio. Each embedding set is modeled as a multivariate Gaussian distribution, $\mathcal{N}_{o}$ for the original and $\mathcal{N}_{r}$ for the reconstructed samples, and FAD is defined as the Fréchet distance between these two distributions.
In addition, we also evaluate our AR model for generating audio using the same metrics. MSE and MAE in this case are calculated considering for each generated sample the closest spectrogram in the NSynth test set.
3.3 Experimental Results
Table 1 compares our MARS with NSynth [10], DDSP [12], and DiffWave [4]. MARS achieves the best scores in NDB/$k$, PKID, and IKID, indicating superior sample diversity and fidelity in both pitch and timbre. It also ranks among the top methods for MSE, MAE, and FAD, reflecting accurate reconstruction and high perceptual quality. PIS and IIS, which measure similarity on sound properties rather than audio quality, remain competitive.
The last row reports metrics on generated audio, which are not directly comparable to the reconstruction metrics; nonetheless, the generated samples maintain low reconstruction error and perceptual similarity, confirming that MARS effectively preserves signal characteristics during synthesis.
3.4 Effectiveness of CMX in Maintaining Audio Quality
In this section, we present a toy example evaluating the impact of CMX. We trained the tokenizer of our AR model for 200k steps using two setups: (i) truncated spectrograms of size $256\times 256\times 1$ obtained from original $512\times 512\times 1$, and (ii) truncated spectrograms rearranged with CMX into $128\times 128\times 4$. Training with CMX reduced training time by $1.5\times$, while still achieving reconstruction quality comparable to the truncated baseline. CMX outperformed truncation in terms of reconstruction accuracy, with lower errors (MSE improved by 0.002 and MAE by 0.022), but showed a slight drop in perceptual quality, as indicated by a 0.19 increase in FAD. Crucially, CMX enables the preservation of all frequency information, which is essential for audio quality, while making this feasible through its substantial memory reduction. Indeed, frequency truncation leads to severe degradation (FAD gap of 2.16 and MSE gap of 0.01 compared to our experiments on input of dimension $256\times 256\times 2$), confirming that CMX is necessary to balance efficiency with fidelity. Overall, these results demonstrate that CMX reduces memory costs, improves reconstruction fidelity, and preserves frequency resolution, with only a marginal trade-off in perceptual quality.
4 Conclusion
In this work, we introduced MARS (Multi-channel AutoRegression on Spectrograms), a novel framework for audio generation that adapts next-scale autoregression from image synthesis to the spectrogram domain. By introducing the channel multiplexing (CMX) technique, MARS reduces spatial resolution while preserving frequency information, enabling scalable training with a manageable memory footprint. A shared tokenizer across scales further ensures consistent discrete representations, allowing the autoregressive model to refine spectrograms hierarchically.
The experimental evaluation on the NSynth dataset demonstrated that MARS achieves competitive or superior performance compared to state-of-the-art baselines across multiple metrics, confirming both the effectiveness of CMX in balancing efficiency with fidelity and the robustness of the overall framework for high-quality audio generation.